OpenStudy (anonymous):

determine all the solutions of the following equations: 1) (cos^2x)(senx) = cos x

OpenStudy (lgbasallote):

do you mean to prove? and im assuming senx is secx LEFT HAND SIDE: \(\Large (\cos^2 x)(sec x) = (\cos^2 x)(\frac{1}{\cos x}) = \cos x = RHS\) rhs -> right hand side

OpenStudy (anonymous):

no, it's not to prove it ... and the equation is with sen not sec

OpenStudy (lgbasallote):

what trig function is sen o.O